OpenStudy (anonymous):

Given \[B=\left[\begin{matrix}2 & 1 \\ 3 & -1\end{matrix}\right]\]Find the matrix of the linear transformation \[g:M _{2}(\mathbb{R})\rightarrow M _{2}(\mathbb{R})\] given by \[g(A)=AB\] with respect to the basis \[E _{11}=\left[\begin{matrix}1 & 0 \\ 0 & 0\end{matrix}\right], E _{12}=\left[\begin{matrix}0 & 1 \\ 0 & 0\end{matrix}\right],E _{21}=\left[\begin{matrix}0 & 0 \\ 1 & 0\end{matrix}\right],E _{22}=\left[\begin{matrix}0 & 0 \\ 0 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

When you have a transformation, and you want the matrix representing that transformation, you need to calculate the transformation of the standard basis "vectors". So what you need to calculate is: \[g(E_{11}), g(E_{12}), g(E_{21}), g(E_{22})\]

OpenStudy (anonymous):

oh okay. does the \[E _{11},E _{12}E _{21},E _{22}\] have any significance? does the 11, 12, 21 and 22 represent anything?

OpenStudy (anonymous):

thats just letting you know where the "1" inside the matrix is. "11" means "first row, first column" "21" means "second row, first column" "ij" means "ith row, jth column"

OpenStudy (anonymous):

oh all right. and does B represent the transformation?

OpenStudy (anonymous):

joe can u do this http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e4f59c60b8b958804a68b7e i am kind of stuck

OpenStudy (anonymous):

The B is pretty much the effect of the transformation. We give the transformation an input, say the matrix A, and what the transformation does is calculates AB.

OpenStudy (anonymous):

oh wait. so you mean that we're transforming all the E matrices into a new matrix by using B? and i have found the g(E) already. not sure why they wrote the matrix in the form of columns :/

OpenStudy (anonymous):

here is what i got for the g(E)s. i worked the first one out in full, then just put the answers for the other 3.

OpenStudy (anonymous):

yeah i got that. but the answer they gave me was to write everything in 1 big matrix where E11 was written in the first column E12 in the 2nd column and so on.

OpenStudy (anonymous):

as columns? thats a little odd. do you mind posting the answer so i can see it?

OpenStudy (anonymous):

yeah sure. hold on a sec.

OpenStudy (anonymous):

alright, i see where they are going now. so after you computed g(E_11) we got: 2 1 0 0 Thats the first column of the answer. the second third and fourth columns are g(E_12), g(E_21), and g(E_22). So they basically created a 4x4 matrix that does the work of the transformation. lets say you had a 2x2 matrix A equal to: a b c d Then you could either compute the transformation by doing g(A) = AB, or by making a 4x1 vector (a, b, c, d) and multiplying it with that 4x4 matrix

OpenStudy (anonymous):

oh. so it is just a different way of expressing it rather than having 4 smaller matrices?

OpenStudy (anonymous):

right. sometimes people dont like dealing with a "vector" space of 2x2 matrices. so instead they can work with the vector space R^4, which is a little more natural to work in.

OpenStudy (anonymous):

Given \[ f(e _{1})=(1,-1,0), f(e _{2})=(0,1,-1),f(e _{3})=(1,0,-1)\]. What is the rank and nullity?

OpenStudy (dumbcow):

If f is a column vector then it yields the following 3X3 matrix in R^3 1 0 1 -1 1 0 0 -1 -1 Add row 2 with row 1 1 0 1 0 1 1 0 -1 -1 Add row3 with row 2 1 0 1 0 1 1 0 0 0 Leaving 2 non-zero rows Thus rank = 2

OpenStudy (dumbcow):

im not sure about nullity?

OpenStudy (dumbcow):

ahh i found it rank + nullity = N since rank = 2, and N=3 nullity = 1

OpenStudy (anonymous):

nice :) i wasnt sure about it either. i dont see that terminology that much.

OpenStudy (dumbcow):

no i did a little research linear algebra was not my fav subject :(